OpenStudy (anonymous):

substitute the given values into the given formula and solve for the unknown variable. If needed be round to 1 decimal place. P=a+b+c;P=40,a=11,b=11 C=??

OpenStudy (anonymous):

\[40/121\]

OpenStudy (anonymous):

\[40=11+11+c\] \[40=22+c\] subtract 22 from both sides \[c= 18\]